(2a-1/2) (2a-1/2). Find solution using suitable identity can u pls tell which identity is used here

Grade:8

1 Answers

Debarati
2 years ago
Here we will use the identity (a-b)2= a2-2ab+b2
(2a-1/2)(2a-1/2)= (2a-1/2)2= (2a)2-(2a)(1/2)+(1/2)2= 4a2-a+1/4= 16a2-4a+1

I hope you understood.